In a saturated solution of Al(OH)₃, the concentration of OH⁻ is 2.6 x 10⁻⁵ mol/L. The concentration of Al³⁺ in mol/ L is which of the following?


  1. 7.8 x 10⁻⁵
  2. 8.7 x 10⁻⁶
  3. 7.8 x 10⁻¹⁵
  4. 2.6 x 10⁻⁵
  5. 5.2 x 10⁻⁵
1
Expert's answer
2022-01-17T17:03:06-0500

Solution:

Al(OH)3(s) ⇌ Al3+(aq) + 3OH(aq)

___S_________S__________3S_____

According to the equation above:

[OH] = 3S = 2.6×10⁻5 mol/L

Hence,

[Al3+] = S = [OH] / 3 = (2.6×10−5 mol/L) / 3 = 8.667×10−6 mol/L = 8.7×10−6 mol/L

[Al3+] = 8.7×106 mol/L


Answer: 2. 8.7×106